a number is divided by 4 and then increased by 160. then 40 is subtracted from the result. the final number is 320 whats the original number?
Answer by checkley79(3341)

X/4+160-40=320
X/4=320-160+40
X/4=200
X=4*200
X=800 ans.
PROOF:
800/4+160-40=320
200+160-40=320
360-40=320
320=320